OpenStudy (anonymous):

(logx)^2 + log(x^3) +2 =0 solve for x

myininaya (myininaya):

so you know we can rewrite log(x^3) as 3log(x) so do you know how to solve u^2+3u=0?

OpenStudy (anonymous):

oh ty let me solve now and get back to you

myininaya (myininaya):

k

OpenStudy (anonymous):

so u(u+3) u=logx so when logx =0 and when log (x)+3 =0 ?

myininaya (myininaya):

right

myininaya (myininaya):

you are good so far

OpenStudy (anonymous):

1 and 10^-3

OpenStudy (anonymous):

I think you forgot the +2 part

myininaya (myininaya):

yeah that is what i got

myininaya (myininaya):

oh i didn't see the plus 2 part

myininaya (myininaya):

@illuminator123 do you see what sour pointed out?

myininaya (myininaya):

we are now looking at \[u^2+3u+2=0\]

OpenStudy (anonymous):

yes u+2 and U+1 are the 2 binomials

myininaya (myininaya):

ok great and you said log(x)=u so you have log(x)+2=0 or log(x)+1=0

OpenStudy (anonymous):

10^-2 10^-1 =x

myininaya (myininaya):

yep yep gj

OpenStudy (anonymous):

thx